# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

### v3

@@ Abstract
The use of copyrighted books for training AI models has led to numerous lawsuits from authors concerned about AI’s ability to generate derivative content. Yet it’s unclear whether these models can generate high quality literary text while emulating authors’ styles/voices. To answer this we conducted a preregistered study comparing MFA-trained expert writers with three frontier AI models: ChatGPT, Claude, and Gemini in writing up to 450 word excerpts emulating 50 award-winning authors’ (including Nobel laureates, Booker Prize winners, and young emerging National Book Award finalists) diverse styles. In blind pairwise evaluations by 159 representative expert (MFA candidates from top U.S. writing programs) and lay readers (recruited via Prolific), AI-generated text from in-context prompting was strongly disfavored by experts for both stylistic fidelity (odds ratio [OR]=0.16, $p<10^{-8}$) and writing quality (OR=0.13, $p<10^{-7}$) but showed mixed results with lay readers. However, fine-tuning ChatGPT on individual author’s complete works completely reversed these findings: experts now favored AI-generated text for stylistic fidelity (OR=8.16, $p<10^{-13}$) and writing quality (OR=1.87, p=0.010), with lay readers showing similar shifts. These effects are robust under cluster-robust inference and generalize across authors and styles in author-level heterogeneity analyses. The fine-tuned outputs were rarely flagged as AI-generated (3% rate versus 97% for in-context prompting) by state-of-the-art AI detectors. Mediation analysis reveals this reversal occurs because fine-tuning eliminates detectable AI stylistic quirks (e.g., cliché density) that penalize in-context outputs, altering the relationship between AI detectability and reader preference. While we do not account for additional costs of human effort required to transform raw AI output into cohesive, publishable novel length prose, the median fine-tuning and inference cost of \$81 per author represents a dramatic 99.7% reduction compared to typical professional writer compensation. Author-specific fine-tuning thus enables non-verbatim AI writing that readers prefer to expert human writing, thereby providing empirical evidence directly relevant to copyright’s fourth fair-use factor, the “effect upon the potential market or value” of the source works.
@@
The U.S. publishing industry supports hundreds of thousands of jobs while generating \$30 billion in yearly revenue, contributing to the larger American copyright sectors that account for \$2.09 trillion in annual GDP contributions (?). Adult Fiction and Non-Fiction books alone accounted for \$6.14 billion in 2024 (?). This economically important sector now faces an unprecedented challenge: its core products have become essential training data for generative-AI systems. Models trained on well-edited books produce more coherent, accurate responses—something crucial to creating the illusion of intelligence (?). Most technology companies building AI use massive datasets of books, typically without permission or licensing (?), and frequently from illegal sources. In the recent copyright lawsuit Bartz vs Anthropic (?), Judge Alsup noted that Anthropic acquired at least five million books from LibGen and two million from Pirate Library Mirror (PiLiMi). Anthropic also used Books3 (a dataset of approximately 191,000 books also used by Meta and Bloomberg to train their language models), now removed after a legal complaint by anti-piracy group, the Rights Alliance. The Bartz vs Anthropic lawsuit also revealed how Anthropic cut millions of print books from their bindings, scanned them into digital files, and threw away the originals solely for the purpose of training Claude. This unauthorized use has sparked outrage among authors (?), triggering dozens of lawsuits against technology companies including OpenAI, Anthropic, Microsoft, Google, and Meta.
@@
While it is widely established that most state-of-the-art Large Language Models (LLMs) have been trained on copyrighted books, it remains unclear whether such training can produce expert-level creative writing. Past research has shown that AI cannot produce high-brow literary fiction or creative nonfiction through prompting alone when compared to professionally trained writers (?). More recent work demonstrates that AI-generated creative writing still remains characterized by clichés, purple prose, and unnecessary exposition  (?). Additionally, relying on Generative AI for creative writing reduces the collective diversity of novel content (?). AI often produces formulaic, mediocre creative writing because it lacks the distinctive personal voice that typically distinguishes one author from another (?). As Pulitzer Fiction finalist Vauhini Vara observes, “ChatGPT’s voice is polite, predictable, inoffensive, upbeat. Great characters, on the other hand, aren’t polite; great plots aren’t predictable; great style isn’t inoffensive; and great endings aren’t upbeat” (?). To address this limitation, practitioners now increasingly prompt AI systems to perform style/voice mimicry by emulating specific writers’ choices (?). This practice has become so common that a fantasy author recently published a novel containing an accidentally included AI prompt requesting emulation of another writer’s style (?). While the effectiveness of such stylistic emulation remains contested, the more pressing question concerns whether style/voice mimicry genuinely improves AI-generated text quality and whether readers—both experts and non-experts—perceive these improvements as meaningful.
@@
We selected closed-source LLMs readily accessible to users without technical expertise: GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. We also tried open-weights Llama 3.1 model but its empirical performance was not good at following long context instructions at the time of our study. Both human experts and LLMs were given the same task: write an excerpt of up to 450 words emulating the style and voice of one of 50 internationally acclaimed authors, including Nobel laureates, Booker Prize winners, and Pulitzer Prize winners, spanning multiple continents and cultures. Among others, our study includes Nobel laureates Han Kang and Annie Ernaux; Booker Prize winners Salman Rushdie, Margaret Atwood, and George Saunders; and Pulitzer Prize winners Junot Díaz and Marilynne Robinson.
Our study examines shorter (up to 450-word) excerpts for two reasons. First, an excerpt of this length isolates the voice, style, and quality that writers must sustain across any longer work. Second, in its current form, an LLM cannot write long-form narratives that automatically balance coherence, quality, thematic consistency, and plot development across thousands of words. But it is merely a matter of time before LLMs reach that capacity. Even now, with human steering and iterative prompting, it is totally possible to produce long-form fiction and non-fiction using shorter model-generated excerpts. This is already happening in real life with startups such as Sudowrite or Genre Fiction publisher Inkitt  (?) specifically focusing on helping consumers write full books using AI. In self-publishing marketplaces such as Kindle, these AI-generated books have already gained popularity  (?, ?).
@@
Figure 1: Figure showing our study design. (1) Select a target author and prompt. (2) Generate upto 450-word candidate excerpts from MFA experts and from LLMs under two settings: in-context prompting (instructions + few-shot examples) and author-specific fine-tuning (model fine-tuned on that author’s works). (3) Readers (experts and lay ) perform blinded, pairwise forced-choice evaluations on two outcomes: stylistic fidelity to the target author and overall writing quality. Pair order and left/right placement are randomized on every trial.
1 Results
@@
Figure 2: (A-B) Forest plots showing odds ratios (OR) and 95% confidence intervals comparing AI and human experts in pairwise evaluations of stylistic fidelity (A) and writing quality (B) where values >>1 favor AI and values <<1 favor humans. Expert readers show preference for human writing when prompted in an in-context setting (OR = 0.16 and 0.13) but that changes when AI is fine-tuned (OR = 8.16 and 1.87). Lay readers have a harder time discriminating, given how they prefer the quality of AI writing even for in-context prompting (OR = 1.55). (C-D) Probability of choosing AI excerpts across individual language models for stylistic fidelity (C) and writing quality (D). Error bars represent 95% confidence intervals. Dashed line indicates chance performance (50%). (E) AI detection accuracy with chosen threshold of $\tau$=0.9 using two state-of-the-art AI detectors (Pangram and GPTZero). Human written text was never misclassified (0.00), in-context AI was detected with 97% accuracy by Pangram and 91% by GPTZero, but fine-tuned AI evaded detection 97% of the time (0.03) for Pangram and 100% of the time in case of GPTZero. (F) Relationship between AI detectability (Pangram) and preference for writing quality across detection score bins. For in-context prompting setup, higher detection scores correlated with lower AI preference (negative slope). This relationship disappeared when AI is fine-tuned (flat slopes). Fine-tuning on an author’s complete oeuvre gets rid of AI quirks while achieving expert-level performance. n = 28 expert readers, 131 lay readers; 3,840 pairwise comparisons with robust clustered standard errors.
1.1 Overall Performance Comparisons
@@
Our final data consist of 3,840 paired-choice tasks, with judgments from 28 MFA experts and 131 lay readers. We fit a logit model for each outcome and condition and include dummies for writer type, reader group, and their interaction. We further employ CR2 cluster robust standard errors (?) clustered at the reader-level to account for within-reader correlation in ratings. Our hypotheses, outcomes, design, and analysis closely follow our OSF pre-registration (SI Sections S4-S8); minor deviations are detailed in SI Section S9.
@@
Figure 2A–B presents the odds ratios, with corresponding predicted probabilities shown in Figure 2C–D. Under in-context prompting, expert readers demonstrated strong preference for human-written text. Odds ratios were 0.16 (95% CI: 0.08–0.29, $p<10^{-8}$) for stylistic fidelity and 0.13 (95% CI: 0.06–0.28, $p<10^{-7}$) for writing quality, indicating six- to eight-fold preferences for human excerpts (Fig. 2A–B). Lay readers showed no significant preference regarding stylistic fidelity (OR = 0.86, 95% CI: 0.62–1.19, $p=0.37$) but favored AI-generated text for writing quality (OR = 1.55, 95% CI: 1.09–2.21, $p=0.014$), selecting AI excerpts in 61% of writing quality trials (Fig. 2C–D). Inter-rater agreement reflected this divergence: expert readers achieved $\kappa=0.58$ for stylistic fidelity and $\kappa=0.41$ for writing quality, while lay readers showed minimal agreement among themselves ($\kappa=0.12$ and $\kappa=0.15$, respectively). Given that lay assessments of literary quality and style reflect inherently diverse tastes, we anticipated lower inter-rater agreement compared to expert evaluators. The writer-type $\times$ reader-type interaction was significant for both outcomes ($\chi^{2}_{(3)}=24.9$, $p=1.6\times 10^{-5}$ for fidelity; $\chi^{2}_{(3)}=37.6$, $p=3.5\times 10^{-8}$ for quality).
@@
Fine-tuning on authors’ complete works reversed these preferences. For expert readers, the odds of selecting the AI excerpt were 8.16 times the odds of selecting the human excerpt for stylistic fidelity (OR =8.16=8.16, 95% CI: 4.69–14.2, $p<10^{-13}$) and 1.87 times the odds for writing quality (OR =1.87=1.87, 95% CI: 1.16–3.02, $p=0.010$). Lay readers showed comparable shifts in their preferences (stylistic fidelity OR =8.29=8.29, 95% CI: 5.15–13.3, $p<10^{-17}$; writing quality OR =2.42=2.42, 95% CI: 1.64–3.57, $p<10^{-5}$). Model-based predicted AI win probabilities converged across groups to about 0.74 for stylistic fidelity and 0.58–0.61 for writing quality (Fig. 2C–D), and the writer-type $\times$ reader-type interaction was no longer significant in the fine-tuned models (both $\chi^{2}_{(1)}<1$, $p>0.40$). Inter-rater agreement among experts increased ($\kappa=0.67$ for writing quality; $\kappa=0.54$ for stylistic fidelity), while agreement among lay readers remained low ($\kappa=0.07$ and $\kappa=0.22$, respectively).
@@
1.2 AI Detection and Stylometric Analysis
@@
GPTZero, another state-of-the-art AI detection, tool showed comparable performance but had higher false-positive-rate, so for our subsequent analyses we stuck to Pangram.
@@
Higher AI-detection scores strongly predicted lower preference rates among expert readers in the in-context prompting condition. For stylistic fidelity, each unit increase in detection score reduced the odds of selecting an excerpt by a factor of 6.3 ($\beta=-1.85\pm 0.29$, $p<10^{-9}$); a similar pattern held for writing quality ($\beta=-2.01\pm 0.33$, $p<10^{-9}$). Fine-tuning largely eliminated this negative relationship between detectability and preference (Pangram $\times$ Setting for style: $\beta=2.56\pm 0.81$, $p=0.002$; for quality: $\beta=2.90\pm 0.88$, $p<0.001$). Two-stage mediation analysis (Fig. 4A) demonstrated that stylometric features, particularly cliché density (See Section S3.2 in SI), mediated 16.4% of the detection effect on preference before fine-tuning but a statistically insignificant $1.3\%$ (95% CI includes zero) afterward, indicating that fine-tuning eliminates rather than masks artificial stylistic signatures.
1.3 Author-Level Performance Heterogeneity
@@
Next, we disaggregated our data to unpack heterogeneity at the level of individual authors. Of 30 fine-tuned author models, 27 exceeded parity for stylistic fidelity (median win rate = 0.74, IQR: 0.63–0.86) and 23 for writing quality (median = 0.58, IQR: 0.54–0.74). Using 95% Jeffreys intervals, fine-tuned models significantly outperformed human writers for 19 authors on stylistic fidelity and 10 authors on writing quality (Fig. 3A). These performance differences showed no systematic relationship with fine-tuning corpus size (Fig. 3B). The fine-tuning premium, i.e., the increase in AI preference rates relative to in-context prompting, ranged from 13.7-13.7 to +70.8+70.8 percentage points for stylistic fidelity (29 of 30 positive) and from 20.8-20.8 to +50.0+50.0 percentage points for writing quality (22 of 30 positive). This “premium” likewise showed no correlation with fine-tuning corpus size (Pearson $r<0.1$ for both outcomes; Fig. 4B).
@@
Figure 4: Fine-tuning substantially reduces stylometric signatures of AI text, improves stylistic fidelity and perceived writing quality over in-context prompting, and substantially cuts costs of producing first draft versus professional writers. (A) Mediation analysis linking stylometric features $\rightarrow$ AI-detector score $\rightarrow$ human preference. Standardized logistic coefficients with 95% CIs are shown for three features for in-context prompting (red) and fine-tuned models (green). Cliché density mediates 16.4% of the detector effect on choice for in-context prompting but only 1.3% for author fine-tuned models; all three features together mediate 25.4% vs -3.2% for in-context prompted and fine-tuned models respectively. (B) “Fine-tuning premium,” defined as $\Delta$ = P(prefer fine-tuned over human) - P(prefer in-context over human), as a function of fine-tuning corpus size. Top: stylistic fidelity; bottom: writing quality. Points are authors; colors denote improvement (green, $\Delta>0$), no change (gray), or degradation (red, $\Delta<0$). Median $\Delta$: +41.7 (fidelity) and +16.7 percentage points (quality). (C) Cost to produce 100,000 words of raw text vs. publishable prose. Expert writers in our study would earn \$25,000 for a 100k-word novel-length manuscript (red). By contrast, AI pipelines can generate 100k words of raw text for \$25–\$276 depending on fine-tuning corpus size (green bars = fine-tuning; hatched = in-context prompting, \$3). This figure reflects direct compute/API costs only, not the additional human steering, chunking, and editing required to turn raw AI text into a cohesive publishable work. Authors ordered by total AI cost.
1.4 Cost Analysis
@@
The weak dependence of performance on fine-tuning corpus size has direct economic implications. Model fine-tuning costs ranged from \$25 to \$276 per author (median = \$81), assuming API-based fine-tuning at \$25 per million tokens plus \$3 for generating 100,000 words of raw text (Fig. 4C). These costs represent approximately 0.3% of what expert writers in our study would charge for a novel-length (100,000 words) manuscript. It should be noted that this comparison reflects raw generation costs before the human steering and editing required to transform AI outputs into publishable works. Despite this caveat, the minimal investment yielded outputs that achieved expert-level performance for most authors. Performance gains were uncorrelated with both corpus size and fine-tuning cost, indicating that computational scale did not drive improvements. The 99.7% reduction in raw generation costs, coupled with superior quality ratings for the majority of authors, underscores the potential for substantial producer-surplus shifts and market displacement.
2 Discussion

### v4

@@ Abstract
The use of copyrighted books for training AI models has led to numerous lawsuits from authors concerned about AI’s ability to generate derivative content. Yet it’s unclear whether these models can generate high quality literary text while emulating authors’ styles/voices. To answer this we conducted a preregistered study comparing MFA-trained expert writers with three frontier AI models: ChatGPT, Claude, and Gemini in writing up to 450-word excerpts emulating 50 award-winning authors’ (including Nobel laureates, Booker Prize winners, and young emerging National Book Award finalists) diverse styles. In blind pairwise evaluations by 28 MFA-trained readers (writers from top U.S. writing programs) and 516 college-educated general readers (recruited via Prolific), AI-generated text from in-context prompting was strongly disfavored by MFA-trained readers for both stylistic fidelity (odds ratio [OR]=0.16) and writing quality (OR=0.13), while college-educated general readers showed no significant preference on stylistic fidelity (OR=1.06) but favored AI-generated text for writing quality (OR=1.82). However, fine-tuning ChatGPT on individual author’s complete works reversed these findings: MFA-trained readers now favored AI-generated text for stylistic fidelity (OR=8.16) and writing quality (OR=1.87), while college-educated general readers showed even stronger AI preference (stylistic fidelity OR=16.65; writing quality OR=5.42). Both reader groups preferred fine-tuned AI, but the writer-type $\times$ reader-type interaction remained significant after fine-tuning ($p=0.021$ for fidelity; $p<10^{-4}$ for quality), indicating that college-educated general readers favored AI-generated text by a wider margin than MFA-trained readers. These effects are robust under cluster-robust inference and generalize across authors and styles in author-level heterogeneity analyses. The fine-tuned outputs were rarely flagged as AI-generated (3% rate versus 97% for in-context prompting) by state-of-the-art AI detectors. Mediation analysis reveals this reversal occurs because fine-tuning eliminates detectable AI stylistic quirks (e.g., cliché density) that penalize in-context outputs, altering the relationship between AI detectability and reader preference. While we do not account for additional costs of human effort required to transform raw AI output into cohesive, publishable novel-length prose, the median fine-tuning and inference cost of \$81 per author represents a dramatic 99.7% reduction compared to typical professional writer compensation. Author-specific fine-tuning thus enables non-verbatim AI writing that readers prefer to expert human writing, thereby providing empirical evidence directly relevant to copyright’s fourth fair-use factor, the “effect upon the potential market or value” of the source works.
@@
The U.S. publishing industry supports hundreds of thousands of jobs while generating \$30 billion in yearly revenue, contributing to the larger American copyright sectors that account for \$2.09 trillion in annual GDP contributions (?). Adult Fiction and Non-Fiction books alone accounted for \$6.14 billion in 2024 (?). This economically important sector now faces an unprecedented challenge: its core products have become essential training data for generative-AI systems. Models trained on well-edited books produce more coherent, accurate responses—something crucial to creating the illusion of intelligence (?). Most technology companies building AI use massive datasets of books, typically without permission or licensing (?), and frequently from illegal sources. In the recent copyright lawsuit Bartz vs Anthropic (?), Judge Alsup noted that Anthropic acquired at least five million books from LibGen and two million from Pirate Library Mirror (PiLiMi). Anthropic also used Books3 (a dataset of approximately 191,000 books also used by Meta and Bloomberg to train their language models), now removed after a legal complaint by anti-piracy group, the Rights Alliance. The Bartz v. Anthropic lawsuit also revealed how Anthropic cut millions of print books from their bindings, scanned them into digital files, and threw away the originals solely for the purpose of training Claude. This unauthorized use has sparked outrage among authors (?), triggering dozens of lawsuits against technology companies including OpenAI, Anthropic, Microsoft, Google, and Meta.
@@
While it is widely established that most state-of-the-art Large Language Models (LLMs) have been trained on copyrighted books, it remains unclear whether such training can produce expert-level creative writing. Past research has shown that AI cannot produce high-brow literary fiction or creative nonfiction through prompting alone when compared to professionally trained writers (?). More recent work demonstrates that AI-generated creative writing still remains characterized by clichés, purple prose, and unnecessary exposition  (?). Additionally, relying on Generative AI for creative writing reduces the collective diversity of novel content (?). AI often produces formulaic, mediocre creative writing because it lacks the distinctive personal voice that typically distinguishes one author from another (?, ?). As Pulitzer Fiction finalist Vauhini Vara observes, “ChatGPT’s voice is polite, predictable, inoffensive, upbeat. Great characters, on the other hand, aren’t polite; great plots aren’t predictable; great style isn’t inoffensive; and great endings aren’t upbeat” (?). To address this limitation, practitioners now increasingly prompt AI systems to perform style/voice mimicry by emulating specific writers’ choices (?). This practice has become so common that a fantasy author recently published a novel containing an accidentally included AI prompt requesting emulation of another writer’s style (?). While the effectiveness of such stylistic emulation remains contested, the more pressing question concerns whether style/voice mimicry genuinely improves AI-generated text quality and whether MFA-trained readers and college-educated general readers perceive these improvements as meaningful.
@@
We selected closed-source LLMs readily accessible to users without technical expertise: GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. We also tried open-weights Llama 3.1 model but its empirical performance was not good at following long context instructions at the time of our study. Both human experts and LLMs were given the same task: write an excerpt of up to 450 words emulating the style and voice of one of 50 internationally acclaimed authors, including Nobel laureates, Booker Prize winners, and Pulitzer Prize winners, spanning multiple continents and cultures. Among others, our study includes Nobel laureates Han Kang and Annie Ernaux; Booker Prize winners Salman Rushdie, Margaret Atwood, and George Saunders; and Pulitzer Prize winners Junot Díaz and Marilynne Robinson. Beyond style emulation, our writing task requires originality under explicit content constraints. Although both AI and human writers received prompts with detailed content specifications, composing literary fiction/creative non-fiction still requires agency especially in deliberate choices of words, syntax, voice, and narrative framing to produce novel, coherent prose. Because literary fiction/creative non-fiction is inherently non-formulaic, the writing task extends beyond style imitation to creative composition, making it different from mere parody or pastiche.
Our study examines shorter (up to 450-word) excerpts for two reasons. First, an excerpt of this length isolates the voice, style, and quality that writers must sustain across any longer work. Second, in its current form, an LLM cannot write long-form narratives that automatically balance coherence, quality, thematic consistency, and plot development across thousands of words. But it is merely a matter of time before LLMs reach that capacity. It should also be noted that human novel-writing is itself iterative and compositional, making excerpt-level quality a meaningful building block. With human steering and iterative prompting, it is feasible to produce long-form fiction and non-fiction using shorter model-generated excerpts. This is already happening in real life with startups such as Sudowrite or Genre Fiction publisher Inkitt  (?) specifically focusing on helping consumers write full books using AI. In self-publishing marketplaces such as Kindle, these AI-generated books have already gained popularity  (?, ?, ?).
@@
Figure 1: Figure showing our study design. (1) Select a target author and prompt. (2) Generate up to 450-word candidate excerpts from MFA experts and from LLMs under two settings: in-context prompting (instructions + few-shot examples) and author-specific fine-tuning (model fine-tuned on that author’s works). (3) Readers (MFA-trained and college-educated) perform blinded, pairwise forced-choice evaluations on two outcomes: stylistic fidelity to the target author and overall writing quality. Pair order and left/right placement are randomized on every trial.
Results
@@
Figure 2: (A-B) Forest plots showing odds ratios (OR) and 95% confidence intervals comparing AI-generated and human-written excerpts in pairwise evaluations of stylistic fidelity (A) and writing quality (B) where values >>1 favor AI and values <<1 favor humans. MFA-trained readers show preference for human writing when prompted in an in-context setting (OR = 0.16 and 0.13) but that changes when AI is fine-tuned (OR = 8.16 and 1.87). College-educated general readers show a distinct preference pattern rather than mere noise: under in-context prompting they are at parity with humans on stylistic fidelity (OR = 1.06) but prefer AI-generated text on writing quality (OR = 1.82). (C-D) Probability of choosing AI excerpts across individual language models for stylistic fidelity (C) and writing quality (D). Error bars represent 95% confidence intervals. Dashed line indicates chance performance (50%). (E) AI detection accuracy with chosen threshold of $\tau$=0.9 using two state-of-the-art AI detectors (Pangram and GPTZero). Human written text was never misclassified (0.00), in-context AI was detected with 97% accuracy by Pangram and 91% by GPTZero, but fine-tuned AI evaded detection 97% of the time (0.03) for Pangram and 100% of the time in case of GPTZero. (F) Relationship between AI detectability (Pangram) and preference for writing quality across detection score bins. For the in-context prompting setup, higher detection scores predict lower AI preference mainly among MFA-trained readers. After fine-tuning, that negative detectability penalty is strongly attenuated, and the model-implied slopes are flat-to-positive. Fine-tuning on an author’s complete oeuvre gets rid of AI quirks while achieving expert-level performance. n = 28 MFA-trained readers, 516 college-educated general readers; 10,920 reader-level pairwise judgments/evaluations.
Overall Performance Comparisons
@@
Our final data consist of 10,920 reader-level pairwise judgments/evaluations, with judgments from 28 MFA-trained readers and 516 college-educated general readers. We fit a logit model for each outcome and condition and include dummies for writer type, reader group, and their interaction. We further employ CR2 cluster robust standard errors (?) clustered at the reader-level to account for within-reader correlation in ratings. Our hypotheses, outcomes, design, and analysis closely follow our OSF pre-registration (SI Sections S4-S8); minor deviations are detailed in SI Section S9.
@@
Figure 2A–B presents the odds ratios, with corresponding predicted probabilities shown in Figure 2C–D. Under in-context prompting, MFA-trained readers demonstrated strong preference for human-written text. Odds ratios were 0.16 (95% CI: 0.08–0.29, $p<10^{-8}$) for stylistic fidelity and 0.13 (95% CI: 0.06–0.28, $p<10^{-6}$) for writing quality, indicating six- to eight-fold preferences for human excerpts (Fig. 2A–B). College-educated general readers showed no significant preference regarding stylistic fidelity (OR = 1.06, 95% CI: 0.90–1.25, $p=0.50$) but clearly favored AI-generated text for writing quality (OR = 1.82, 95% CI: 1.53–2.15, $p<10^{-10}$), selecting AI excerpts in 50–62% of writing quality trials depending on the model (Fig. 2C–D). Inter-rater agreement reflected this divergence: MFA-trained readers achieved $\kappa=0.58$ for stylistic fidelity and $\kappa=0.41$ for writing quality, while college-educated readers showed minimal agreement among themselves ($\kappa=0.10$ and $\kappa=0.08$, respectively). Given that college-educated general readers’ assessments of literary quality and style reflect inherently diverse tastes, we anticipated lower inter-rater agreement compared to MFA-trained evaluators. The writer-type $\times$ reader-type interaction was significant for both outcomes ($\chi^{2}_{(3)}=36.6$, $p=5.6\times 10^{-8}$ for fidelity; $\chi^{2}_{(3)}=46.9$, $p=3.7\times 10^{-10}$ for quality).
@@
Fine-tuning on authors’ complete works reversed these preferences. For MFA-trained readers, the odds of selecting the AI excerpt were 8.16 times the odds of selecting the human excerpt for stylistic fidelity (OR =8.16=8.16, 95% CI: 4.69–14.2, $p<10^{-12}$) and 1.87 times the odds for writing quality (OR =1.87=1.87, 95% CI: 1.16–3.02, $p=0.010$). College-educated general readers showed even larger shifts in their preferences (stylistic fidelity OR = 16.65, 95% CI: 13.03–21.28, $p<10^{-15}$; writing quality OR = 5.42, 95% CI: 4.41–6.66, $p<10^{-15}$). Unlike the in-context condition, both reader groups now favored fine-tuned AI, but the magnitude of this preference differed: model-based predicted AI win probabilities were 0.74 for MFA-trained readers and 0.80 for college-educated general readers on stylistic fidelity, and 0.58 versus 0.70 on writing quality (Fig. 2C–D). The writer-type $\times$ reader-type interaction remained significant in the fine-tuned models (stylistic fidelity $\chi^{2}_{(1)}=5.33$, $p=0.021$; writing quality $\chi^{2}_{(1)}=15.98$, $p=6.4\times 10^{-5}$), indicating that college-educated general readers favored AI-generated text by a wider margin than MFA-trained readers even after fine-tuning. Inter-rater agreement among MFA-trained readers increased ($\kappa=0.67$ for writing quality; $\kappa=0.54$ for stylistic fidelity), while agreement among college-educated general readers remained low ($\kappa=0.12$ and $\kappa=0.26$, respectively).
@@
AI Detection and Stylometric Analysis
@@
GPTZero, another state-of-the-art AI detection, tool showed comparable performance but had higher false-positive rate, so we used Pangram for our subsequent analyses.
@@
Higher AI-detection scores strongly predicted lower preference rates among MFA-trained readers in the in-context prompting condition. For stylistic fidelity, each unit increase in detection score reduced the odds of selecting an excerpt by a factor of 6.3 ($\beta=-1.85\pm 0.29$, $p<10^{-9}$); a similar pattern held for writing quality ($\beta=-2.01\pm 0.33$, $p<10^{-8}$). College-educated general readers were substantially less sensitive to detectability even before fine-tuning (Pangram $\times$ Reader Type for style: $\beta=1.88\pm 0.30$, $p<10^{-9}$; for quality: $\beta=2.61\pm 0.34$, $p<10^{-13}$). Fine-tuning attenuated the negative relationship between detectability and preference (Pangram $\times$ Setting for style: $\beta=2.56\pm 0.81$, $p=0.002$; for quality: $\beta=2.90\pm 0.88$, $p<0.001$), though the degree of attenuation differed across reader groups, as indicated by significant three-way interactions (Pangram $\times$ Setting $\times$ Reader Type for style: $\beta=-1.77\pm 0.84$, $p=0.036$; for quality: $\beta=-3.26\pm 0.91$, $p<0.001$). Two-stage mediation analysis (Fig. 4A) demonstrated that among MFA-trained readers, where the detectability penalty is concentrated, stylometric features, particularly cliché density (See Section S3.2 in SI), mediated 16.4% of the detection effect on preference before fine-tuning but a statistically insignificant 1.3% (95% CI includes zero) afterward, indicating that fine-tuning reduces rather than merely masks artificial stylistic signatures. Because college-educated general readers did not penalize AI-detectable text in either condition, mediation analysis does not apply to this group.
Author-Level Performance Heterogeneity
@@
Next, we disaggregated our data to unpack heterogeneity at the level of individual authors. Of 30 fine-tuned author models, 29 exceeded parity for stylistic fidelity (median win rate = 0.83, IQR: 0.70–0.94) and 28 for writing quality (median = 0.67, IQR: 0.58–0.77). Using 95% Jeffreys intervals, fine-tuned models significantly outperformed human writers for 26 authors on stylistic fidelity and 21 authors on writing quality (Fig. 3A). These performance differences showed no systematic relationship with fine-tuning corpus size (Fig. 3B). The fine-tuning premium, i.e., the increase in AI preference rates relative to in-context prompting, ranged from 12.2-12.2 to +67.4+67.4 percentage points for stylistic fidelity (29 of 30 positive) and from 14.1-14.1 to +46.1+46.1 percentage points for writing quality (24 of 30 positive). This “premium” likewise showed no correlation with fine-tuning corpus size (Pearson $r<0.1$ for both outcomes; Fig. 4B).
@@
Figure 4: Fine-tuning substantially reduces stylometric signatures of AI text, improves stylistic fidelity and perceived writing quality over in-context prompting, and substantially cuts costs of producing first draft versus professional writers. (A) Mediation analysis linking stylometric features $\rightarrow$ AI-detector score $\rightarrow$ human preference. Standardized logistic coefficients with 95% CIs are shown for three features for in-context prompting (red) and fine-tuned models (green). Cliché density mediates 16.4% of the detector effect on choice for in-context prompting but only 1.3% for author fine-tuned models; all three features together mediate 25.4% vs -3.2% for in-context prompted and fine-tuned models respectively. (B) “Fine-tuning premium,” defined as $\Delta$ = P(prefer fine-tuned over human) - P(prefer in-context over human), as a function of fine-tuning corpus size. Top: stylistic fidelity; bottom: writing quality. Points are authors; colors denote improvement (green, $\Delta>0$), no change (gray), or degradation (red, $\Delta<0$). Median $\Delta$: +34.2+34.2 (fidelity) and +13.4+13.4 percentage points (quality). (C) Cost to produce 100,000 words of raw text vs. publishable prose. MFA-trained writers in our study would earn \$25,000 for a 100k-word novel-length manuscript (red). By contrast, AI pipelines can generate 100k words of raw text for \$25–\$276 depending on fine-tuning corpus size (green bars = fine-tuning; hatched = in-context prompting, \$3). This figure reflects direct compute/API costs only, not the additional human steering, chunking, and editing required to turn raw AI text into a cohesive publishable work. Authors ordered by total AI cost.
Cost Analysis
@@
The weak dependence of performance on fine-tuning corpus size has direct economic implications. Total AI generations costs (fine-tuning plus inference) ranged from \$25 to \$276 per author (median = \$81), assuming API-based fine-tuning at \$25 per million tokens plus \$3 for generating 100,000 words of raw text (Fig. 4C). These costs represent approximately 0.3% of what MFA-trained writers in our study would charge for a novel-length (100,000 words) manuscript. It should be noted that this comparison reflects raw generation costs before the human steering and editing required to transform AI outputs into publishable works. Despite this caveat, the minimal investment yielded outputs that achieved expert-level performance for most authors. Performance gains were uncorrelated with both corpus size and fine-tuning cost, indicating that computational scale did not drive improvements. The 99.7% reduction in raw generation costs, coupled with superior quality ratings for the majority of authors, underscores the potential for substantial producer-surplus shifts and market displacement.
Discussion